Hard at work early in the morning? Take a break 
Black - 20/30
Working - 12/20
Blue - 10/30
Working - 4/10
I'm assuming picking 2 pens ONE AFTER ANOTHER, if it's AT THE
SAME TIME then make the necessary changes on your own (which I
guess, is doable)
1st Black and working - 20/30 x 12/20 = 2/5
2nd Black and working - 19/29 x 11/19 = 11/29
Black and working = 2/5 x 11/29 = 22/145
Blue and working = 10/30 x 4/10 x 9/29 x 3/9 = 2/145
Probability = 22/145 + 2/145 = 24/145
Haven't done probability in a long time. Better get someone to
re-confirm first.